OpenStudy (anonymous):

find the lengths of both circular arcs of the unit circle connecting the point (square root 2/2, - square root 2/2) and the point who's radius makes an angle of .97 radians with the positive horizontal axis?

OpenStudy (hoblos):

x=√2/2 y=-√2/2 so the angle by the first point's radius with the positive horizontal axis is -(pi)/4 = -0.785 rad the lengths of the first circular arc= 0.97 + 0.785 = 1.755rad the lengths of the second circular arc= 2(pi)- 1.755 = 5.525rad

OpenStudy (anonymous):

1.755 worked, 5.525 didn't.

OpenStudy (hoblos):

ops sorry sorry typo!! it is 4.525

OpenStudy (anonymous):

alright haha, ty. that did. thanks again!

OpenStudy (hoblos):

0k good :) yw XD